Cullumber Company, organized in 2019, has set up a single account for all intangible assets. The following summary discloses the debit entries that have been recorded during 2020.

1/2/20 Purchased patent (8-year life)

$354,400

4/1/20 Purchase goodwill (indefinite life)

353,000

7/1/20 Purchased franchise with 10-year life; expiration date 7/1/30

433,000

8/1/20 Payment of copyright (5-year life)

159,600

9/1/20 Research and development costs

223,000

$1,523,000

Prepare the necessary entry to clear the Intangible Assets account and to set up separate accounts for distinct types of intangibles. (Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts.)

Account Titles and Explanation

Debit

Credit

enter an account title

enter a debit amount

enter a credit amount

enter an account title

enter a debit amount

enter a credit amount

enter an account title

enter a debit amount

enter a credit amount

enter an account title

enter a debit amount

enter a credit amount

enter an account title

enter a debit amount

enter a credit amount

enter an account title

enter a debit amount

enter a credit amount

Make the entry as of December 31, 2020, recording any necessary amortization. (Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts.)

Account Titles and Explanation

Debit

Credit

enter an account title

enter a debit amount

enter a credit amount

enter an account title

enter a debit amount

enter a credit amount

enter an account title

enter a debit amount

enter a credit amount

enter an account title

enter a debit amount

enter a credit amount

Reflect all balances accurately as of December 31, 2020.

Use straight-line amortization

select an intangible asset                                                                      CopyrightsResearch and development costsPatentsFranchisesGoodwill

$enter a dollar amount

select an intangible asset                                                                      CopyrightsGoodwillFranchisesResearch and development costsPatents

$enter a dollar amount

select an intangible asset                                                                      PatentsResearch and development costsFranchisesGoodwillCopyrights

$enter a dollar amount

select an intangible asset                                                                      GoodwillFranchisesPatentsCopyrightsResearch and development costs

$enter a dollar amount
Answer #1
a
Account Titles and Explanation Debit Credit
Patents 354400
Goodwill 353000
Franchise 433000
Copyright 159600
Research and Development Expenses 223000
       Intangible Assets 1523000
b
Account Titles and Explanation Debit Credit
Amortization Expense 79250
      Patents 44300 =354400/8
      Franchise 21650 =433000/10*6/12
      Copyright 13300 =159600/5*5/12
c
Patents 310100 =354400-44300
Goodwill 353000
Franchises 411350 =433000-21650
Copyrights 146300 =159600-13300